OpenStudy (anonymous):

I need some help with 'Verifying Trigonometric Identities': 13. (sinx + tanx) / (1 + cosx) = tanx 15. cosx = (1 - tan(x/2)^2)/(1 tan(x/2)^2)

OpenStudy (freckles):

\[\frac{\sin(x)+\frac{\sin(x)}{\cos(x)}}{1+\cos(x)}\]

OpenStudy (freckles):

multiply top and bottom by cos(x) to get rid of that compound fraction as a first step

OpenStudy (anonymous):

\[\frac{ \sin \theta \cos \theta + \sin \theta}{ \cos \theta + \cos ^{2} \theta }\] This is what I come up with, I'm not sure what to do from here..

OpenStudy (freckles):

Ok good stuff I wouldn't have multiplied the bottom all out but that is fine... We are going to factor sin(x) on top and cos(x) on bottom you should see something that cancels

OpenStudy (freckles):

\[\frac{\sin(x)(\cos(x)+1)}{\cos(x)(1+\cos(x))}\]

OpenStudy (anonymous):

Awesome, thanks! Any chance you could point me in the right direction for the second one?

OpenStudy (freckles):

I'm assuming there is a type-o in the second one

OpenStudy (freckles):

Is that a wrong assumption?

OpenStudy (anonymous):

\[\cos x = \frac{ 1 - \tan ^{2} \frac{ x }{ 2 } }{ 1 + \tan ^{2} \frac{ x }{ 2 }}\]

OpenStudy (freckles):

I usually prefer things in terms of sin and cos so I don't have to remember extra identities

OpenStudy (freckles):

but if you know tan(x/2) 's identity we can use that here

OpenStudy (freckles):

do you know the half angle identity for tan?

OpenStudy (freckles):

It is fine if you don't like I said I prefer writing things in terms of sin and cos

OpenStudy (anonymous):

\[\frac{ \sin \alpha }{ 1 + \cos \alpha } = \tan \frac{ \alpha }{ 2 }\]

OpenStudy (anonymous):

The upcoming test is over all of the trigonometric identities so it's best if I use them all.

OpenStudy (anonymous):

Yaknow?

OpenStudy (freckles):

let's do this way and let's see how hard it is the other way this way might actually be easier... \[\frac{1-\frac{\sin^2(\frac{x}{2})}{\cos^2(\frac{x}{2})}}{1+\frac{\sin^2(\frac{x}{2})}{\cos^2(\frac{x}{2})}}\]

OpenStudy (freckles):

try multiplying top and bottom by cos^2(x/2) to get rid of the compound fraction first

OpenStudy (anonymous):

Alright, I got this: \[\frac{ \cos ^{2} \frac{ x }{ 2 } - \sin ^{2}\frac{ x }{ 2 }}{ \cos ^{2} \frac{ x }{ 2 } + \sin ^{2}\frac{ x }{ 2 } }\]

OpenStudy (freckles):

what does the bottom equal ?

OpenStudy (anonymous):

It's equal to 1

OpenStudy (freckles):

yes \[\sin^2(\theta)+\cos^2(\theta)=1 \]

OpenStudy (freckles):

do you know what the top equals?

OpenStudy (anonymous):

I don't think so

OpenStudy (freckles):

\[\cos^2(\theta)-\sin^2(\theta)=\cos(2 \theta) \\ \text{ so if } \theta=\frac{x}{2} \text{ we have } \\ \cos^2(\frac{x}{2})-\sin^2(\frac{x}{2})=\cos(2 \cdot \frac{x}{2})=\cos(x)\]

OpenStudy (freckles):

I will show you the other way if you like now

OpenStudy (freckles):

or if you have a question about this way let me know

OpenStudy (anonymous):

Dude, if you're not too busy, that would be awesome. Either way you've already helped me a ton!

OpenStudy (freckles):

\[\frac{1-\tan^2(\frac{x}{2})}{1+\tan^2(\frac{x}{2})} =\frac{1-(\frac{\sin(x)}{1+\cos(x)})^2}{1+(\frac{\sin(x)}{1+\cos(x)})^2} \\ =\frac{(1+\cos(x))^2}{(1+\cos(x))^2} \cdot \frac{1-(\frac{\sin(x)}{1+\cos(x)})^2}{1+(\frac{\sin(x)}{1+\cos(x)})^2} =\frac{(1+\cos(x))^2-\sin^2(x)}{(1+\cos(x))^2+\sin^2(x)}\] While you look at that I will typing more

OpenStudy (freckles):

\[\frac{1+2\cos(x)+\cos^2(x)-\sin^2(x)}{1+2\cos(x)+\cos^2(x)+\sin^2(x)}=\frac{1+2\cos(x)+\cos^2(x)-(1-\cos^2(x))}{1+2\cos(x)+1}\]

OpenStudy (freckles):

\[\frac{1+2\cos(x)+\cos^2(x)-1+\cos^2(x)}{2+2\cos(x)}=\frac{2\cos(x)+2\cos^2(x)}{2+\cos(x)} \\ =\frac{\cos(x)(2+\cos(x))}{2+\cos(x)}=\cos(x)\]

OpenStudy (freckles):

so a little more longer but still works

OpenStudy (anonymous):

Wow...Thank you so much!

OpenStudy (freckles):

np

OpenStudy (freckles):

good luck on your test

OpenStudy (anonymous):

Thank you~!